Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is.
14

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=14
⇒ m =14/2
= 7
m2–1= 72–1
= 49–1
= 48
m2+1=72+1
= 49+1
=50
∴ (14, 48, 50) is not a Pythagorean triplet.

Thanks